# Exact pp-wave solutions in shift-symmetric higher-order scalar-tensor theories

### 19 Mar 2026 v1

@@ IV Disformal and conformal transformations of the stealth pp-wave solutions
Quadratic-order HOST theories are closed under such transformations. Applying (25) to a theory in this class generates another theory of the same class. Hence the transformation acts as a map between different quadratic-order HOST theories without spoiling their structure.
@@
which remains finite under the same non-degeneracy condition.
Applying (25) to the stealth pp-wave solution (19), (20), and (21), the transformed metric takes the form
@@
Thus the disformal transformation maps a stealth pp-wave solution into another solution with the same Brinkmann form but with a shifted profile $\bar{H}$ and a rescaled kinetic term $\bar{X}$. However, the transformed profile does not in general satisfy the pp-wave equation. Indeed, one finds
@@
The construction naturally extends to the combined conformal–disformal transformation
@@
Under the general conformal–disformal map (36), the same qualitative picture persists: the pp-wave ansatz remains valid, with its amplitude $\bar{H}$ rescaled by $A(X_{0})$ and shifted by $B(X_{0})$. Thus, the family of pp-wave metrics is kinematically stable under both conformal and disformal redefinitions, although the dynamical decoupling between the scalar and tensor sectors is generically lost. Disformal transformations therefore act as a bridge between stealth and non-stealth configurations within the broader HOST framework.
@@
Because $A_{0}$ is constant, the transformation does not distort the null structure of the spacetime: the vector $\partial_{v}$ remains null, the surfaces $u=\text{const.}$ remain null hypersurfaces, and the causal structure is unchanged. In particular, the pp-wave condition
@@
which remains constant and finite for $A_{0}>0$. Even if $A(X)$ has a nontrivial functional dependence on $X$, the constancy of $X$ on the pp-wave background ensures that $A(X_{0})$ acts as a constant. Thus, the pure conformal map does not introduce any new structure: it simply rescales all lengths uniformly. This behavior stands in sharp contrast to the disformal case. A disformal transformation introduces an anisotropic deformation along the direction of $\phi_{\mu}$, thereby modifying the transverse geometry and coupling the scalar and tensor modes. By contrast, a pure conformal transformation preserves the isotropy of the $(x_{1},x_{2})$ plane and does not generate any mixing between the scalar and gravitational sectors. As a result, stealth pp-wave solutions remain stealth under conformal transformations, and hence the scalar field continues to have no backreaction on the geometry, and the propagation of the gravitational wave is unchanged up to an overall rescaling of its amplitude. From a physical perspective, this reflects the fact that a constant conformal rescaling does not alter the curvature tensors in any essential way. The Weyl tensor, which encodes the radiative degrees of freedom of the gravitational field, is invariant under constant conformal transformations, and the Ricci tensor rescales trivially. Therefore, the conformally transformed metric describes the same gravitational wave solution, with the same profile $H(u,x_{i})$ and the same decoupled scalar configuration.
@@
Thus, while the conformal sector acts trivially on pp-wave geometries, the disformal sector generates genuinely new non-stealth configurations within the same HOST class.
@@ V Conclusion
In this work, we have presented a comprehensive analysis of exact plane-fronted gravitational wave (pp-wave) solutions in shift-symmetric quadratic-order higher-order scalar–tensor (HOST) theories, with particular emphasis on the degenerate subclasses (DHOST). By imposing only minimal geometric assumptions— the existence of a null coordinate and a two-dimensional Euclidean transverse space—we constructed the most general pp-wave ansatz obtainable after exhausting the residual coordinate freedom. Substituting this ansatz into the HOST action, we showed that the Euler–Lagrange equations reduce to a single two-dimensional Laplace equation for the metric profile $H(u,x_{1},x_{2})$, exactly as in vacuum general relativity (GR). This demonstrates that the mathematical structure governing pp-waves is remarkably stable under higher-order scalar–tensor extensions.
@@
Overall, our results show that pp-wave geometries provide a robust probe of the structure of higher-order scalar–tensor gravity. They reveal which aspects of GR persist unchanged in DHOST theories, which features are modified, and how scalar degrees of freedom can remain hidden or become dynamically relevant depending on the frame. Future work may explore the phenomenological implications of these solutions for gravitational-wave observations, their generalization to broader HOST models, and the systematic use of disformal transformations as a solution-generating mechanism in scalar–tensor gravity.
@@
Another important direction concerns the observational implications of these solutions. Since pp-waves represent exact gravitational-wave geometries, they provide a controlled setting in which to study how scalar degrees of freedom propagate alongside gravitational radiation. Disformal transformations, in particular, generate configurations in which scalar and tensor waves are intrinsically coupled. Exploring how such couplings manifest in waveform distortions, polarization content, or propagation speed could offer new avenues for constraining scalar–tensor gravity with current and future gravitational-wave detectors.
It would also be valuable to extend the present analysis to more general classes of HOST theories, including models with broken shift symmetry, nontrivial scalar potentials, or higher-order derivative interactions. Whether the Laplace structure of the pp-wave equation persists in these broader settings remains an open question. Similarly, generalizations to other exact wave geometries—such as Kundt waves, gyratons, or impulsive wave solutions—may reveal additional universal features shared between GR and higher-order scalar–tensor theories.
Finally, the role of disformal transformations as a solution-generating mechanism deserves further exploration. The fact that stealth pp-waves are mapped to non-stealth configurations suggests that disformal maps can be used to systematically construct new families of exact solutions with controlled scalar– tensor interactions. Developing a classification of pp-wave–like geometries under the full conformal–disformal group may shed light on the deeper geometric structure underlying DHOST theories and their relation to GR.
@@ A.3 Removal of the $du\,dx^{i}$ terms and gyratonic effects
It is important to emphasize that this removal is generally local. In a global context, if the transverse space has a non-trivial topology or if one considers the angular momentum of the wave source, these $g_{ui}$ terms may not be entirely removable and can encode gyratonic effects [36, 37]. In the present work, we restrict our attention to the case where such terms are globally absent, focusing on the standard pp-wave class without gyratonic contributions.
@@ A.4 Conformal flatness and the Brinkmann form
Finally, we comment on the reduction to the canonical pp-wave form (19). While the steps above are based on coordinate freedom, the additional requirements $I=0$ and $W=1$ are essential to satisfy the geometric conditions $\nabla_{\nu}\ell^{\mu}=0$ and $\ell^{\mu}\ell_{\mu}=0$ as defined in Eq. (7). Without $I=0$, the vector $\ell^{\mu}$ would not be null. And without $W=1$, the null rays would not be parallel in the sense of a plane-fronted wave, and the metric would belong to the more general Kundt class. Thus, the specialization to Eq. (19) is a choice of the specific geometric class of spacetimes being investigated in the HOST analysis.

### 31 Mar 2026 v2

@@ IV Disformal and conformal transformations of the stealth pp-wave solutions
Quadratic-order HOST theories are closed under such transformations. Applying (25) to a theory in this class generates another theory of the same class; hence the transformation acts as a map between different quadratic-order HOST theories without spoiling their structure.
@@
which remains finite under the same nondegeneracy condition.
Applying (25) to the stealth pp-wave solution (19)-(21), the transformed metric takes the form
@@
Thus the disformal transformation maps a stealth pp-wave solution into another solution with the same Brinkmann form but with a shifted profile $\bar{H}$ and a rescaled kinetic term $\bar{X}$. However, the transformed profile does not, in general, satisfy the pp-wave equation. Indeed, one finds
@@
The construction naturally extends to the combined conformal-disformal transformation
@@
Under the general conformal-disformal map (36), the same qualitative picture persists: the pp-wave ansatz remains valid, with its amplitude $\bar{H}$ rescaled by $A(X_{0})$ and shifted by $B(X_{0})$. Thus, the family of pp-wave metrics is kinematically stable under both conformal and disformal redefinitions, although the dynamical decoupling between the scalar and tensor sectors is generically lost. Disformal transformations therefore act as a bridge between stealth and nonstealth configurations within the broader HOST framework.
@@
Because $A_{0}$ is constant, the transformation does not distort the null structure of the spacetime: the vector $\partial_{v}$ remains null, the surfaces $u=\text{const}$ remain null hypersurfaces, and the causal structure is unchanged. In particular, the pp-wave condition
@@
which remains constant and finite for $A_{0}>0$. Even if $A(X)$ has a nontrivial functional dependence on $X$, the constancy of $X$ on the pp-wave background ensures that $A(X_{0})$ acts as a constant. Thus, the pure conformal map does not introduce any new structure: it simply rescales all lengths uniformly. This behavior stands in sharp contrast to the disformal case. A disformal transformation introduces an anisotropic deformation along the direction of $\phi_{\mu}$, thereby modifying the transverse geometry and coupling the scalar and tensor modes. By contrast, a pure conformal transformation preserves the isotropy of the $(x_{1},x_{2})$ plane and does not generate any mixing between the scalar and gravitational sectors. As a result, stealth pp-wave solutions remain stealth under conformal transformations; hence, the scalar field continues to have no backreaction on the geometry, and the propagation of the gravitational wave is unchanged up to an overall rescaling of its amplitude. From a physical perspective, this reflects the fact that a constant conformal rescaling does not alter the curvature tensors in any essential way. The Weyl tensor, which encodes the radiative degrees of freedom of the gravitational field, is invariant under constant conformal transformations, and the Ricci tensor rescales trivially. Therefore, the conformally transformed metric describes the same gravitational-wave solution, with the same profile $H(u,x_{i})$ and the same decoupled scalar configuration.
@@
Thus, while the conformal sector acts trivially on pp-wave geometries, the disformal sector generates genuinely new nonstealth configurations within the same HOST class.
@@ V Conclusion
In this work, we have presented a comprehensive analysis of exact plane-fronted gravitational-wave (pp-wave) solutions in shift-symmetric quadratic-order HOST theories, with particular emphasis on the degenerate subclasses (DHOST). By imposing only minimal geometric assumptions- the existence of a null coordinate and a two-dimensional Euclidean transverse space—we constructed the most general pp-wave ansatz obtainable after exhausting the residual coordinate freedom. Substituting this ansatz into the HOST action, we showed that the Euler-Lagrange equations reduce to a single two-dimensional Laplace equation for the metric profile $H(u,x_{1},x_{2})$, exactly as in vacuum GR. This demonstrates that the mathematical structure governing pp-waves is remarkably stable under higher-order scalar-tensor extensions.
@@
Overall, our results show that pp-wave geometries provide a robust probe of the structure of higher-order scalar-tensor gravity. They reveal which aspects of GR persist unchanged in DHOST theories, which features are modified, and how scalar degrees of freedom can remain hidden or become dynamically relevant depending on the frame. Future work may explore the phenomenological implications of these solutions for gravitational-wave observations, their generalization to broader HOST models, and the systematic use of disformal transformations as a solution-generating mechanism in scalar-tensor gravity.
@@
Another important direction concerns the observational implications of these solutions. Since pp-waves represent exact gravitational-wave geometries, they provide a controlled setting in which to study how scalar degrees of freedom propagate alongside gravitational radiation. Disformal transformations, in particular, generate configurations in which scalar and tensor waves are intrinsically coupled. Exploring how such couplings manifest in waveform distortions, polarization content, or propagation speed could offer new avenues for constraining scalar-tensor gravity with current and future gravitational-wave detectors.
It would also be valuable to extend the present analysis to more general classes of HOST theories, including models with broken shift symmetry, nontrivial scalar potentials, or higher-order derivative interactions. Whether the Laplace structure of the pp-wave equation persists in these broader settings remains an open question. Similarly, generalizations to other exact wave geometries—such as Kundt waves, gyratons, or impulsive wave solutions—may reveal additional universal features shared between GR and higher-order scalar-tensor theories.
Finally, the role of disformal transformations as a solution-generating mechanism deserves further exploration. The fact that stealth pp-waves are mapped to nonstealth configurations suggests that disformal maps can be used to systematically construct new families of exact solutions with controlled scalar- tensor interactions. Developing a classification of pp-wave-like geometries under the full conformal-disformal group may shed light on the deeper geometric structure underlying DHOST theories and their relation to GR.
@@ A.3 Removal of the $du\,dx^{i}$ terms and gyratonic effects
It is important to emphasize that this removal is generally local. In a global context, if the transverse space has a nontrivial topology or if one considers the angular momentum of the wave source, these $g_{ui}$ terms may not be entirely removable and can encode gyratonic effects [39, 40]. In the present work, we restrict our attention to the case where such terms are globally absent, focusing on the standard pp-wave class without gyratonic contributions.
@@ A.4 Conformal flatness and the Brinkmann form
Finally, we comment on the reduction to the canonical pp-wave form (19). While the steps above are based on coordinate freedom, the additional requirements $I=0$ and $W=1$ are essential to satisfy the geometric conditions $\nabla_{\nu}\ell^{\mu}=0$ and $\ell^{\mu}\ell_{\mu}=0$ as defined in Eq. (7). Without $I=0$, the vector $\ell^{\mu}$ would not be null; without $W=1$, the null rays would not be parallel in the sense of a plane-fronted wave, and the metric would belong to the more general Kundt class. Thus, the specialization to Eq. (19) is a choice of the specific geometric class of spacetimes being investigated in the HOST analysis.